### v2

@@ 3.2 Continuum emission
This is evidenced from Fig. 6, where we compare, in the right panels, the predicted in-flight annihilation emission for DM masses of to MeV, with SPI and COMPTEL data in the region of the sky, for a Gondolo-Silk (top row) and a *MIN (bottom row) DM spike benchmarks. For each DM mass, we perform a fit to the keV longitudinal profile (shown in the left panels) and normalize the value to the best-fit, including the astrophysical disk component from stars. The expected IC background emission is shown as a solid black line. For completeness, we show in Table 1 our best-fit values used to obtain the estimations.
From this figure, one can clearly see that masses of up to MeV seem to be well compatible with the current measurements of the diffuse -ray galactic flux, while higher masses will significantly exceed the data (especially, COMPTEL observations) and would exceed the background emission, which is not expected. A similar figure to this one is shown in Appendix A (Figure 12) but for the case of a NFW DM distribution without spike.
5 MeV | 10 MeV | 20 MeV | 50 MeV | 100 MeV | |
@@
A possible test of our proposal would come from data in ROIs other than -, as well as from different experiments. To the best of our knowledge, these are provided by COMPTEL and EGRET, which we show and discuss in App. C, finding that they do not alter our conclusions.
@@ 4 Summary and discussion
For the first time in this context, we have considered the possibility that the DM density profile develops a spike around Sgr A*. This is expected for peaked DM profiles like NFW, that were already known to be needed to possibly explain this signal. We have computed the associated 511 keV line emission and demonstrated that it can fit the bulge data whilst producing very little emission at the disk, which we have assumed to have originated mainly from astrophysical sources, e.g. young stars. Remarkably, this conclusion holds not only for Gondolo-Silk spikes, but also for more conservative ones that take into account DM profile softening by stars in the vicinity of Sgr A* and are supported by simulations (all our spike benchmarks respect the limits on the mass close to Sgr A* by the GRAVITY experiment [37]). DM profiles like NFW without a spike, that explained the observed 511 keV bulge emission, have been recently found to produce too much disk emission –which we confirmed here– and hence they were not favoured by data. Our findings, therefore, motivate new interest in the DM hypothesis for the 511 keV line.
@@
Finally concerning our new results, we have tested how some variations, either in the diffusion setup or in the electron density distribution in the Milky Way, can affect the reported emission profiles. As we discuss in Appendix A, variations of our benchmark scenario mainly lead to a more peaked profile, which leads to an even higher ratio for the bulge-to-disk emission and thus reinforce our conclusions. Furthermore, an inhomogeneous diffusion scenario where CRs are more confined around the GC [82], as motivated by observations of the Fermi-LAT [83, 84, 85], will also lead to a more peaked profile of the predicted line. However, winds in the GC will have the opposite effect.
We now move to comment on observational implications of our results for MeV telescopes. First, we note that the SPI measurements employed are significantly affected by the templates used to extract the observations. In particular, the data that we are using are extracted using Gaussian templates, which have a similar symmetry as the one expected by DM-like models. Recently, stellar templates for the nuclear stellar bulge and a boxy bulge were shown to be statistically preferred over the profile expected from an NFW DM distribution [9], even though positron propagation was not included on top of NFW and thus, strictly speaking, the conclusion applies only to DM masses closest to an MeV. In addition, Ref. [9] found that the NFW template was the best model when analyzing the keV line, while it is the continuum o-ps emission what makes their analysis favour the stellar bulge hypothesis. Moreover, the templates of [9] did not consider the DM spike profiles that we are exploring here. Therefore, our goal here is not to provide the best and most accurate explanation of the keV line emission, what would require to have access to SPI’s raw data and the coded mask response function, but to demonstrate that the keV observations and the correlated in-flight annihilation emission motivate considering the possibility of an origin from DM annihilations alleviating previous constraints.
@@ Acknowledgements
SB and MF are supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union “NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, wich are both funded “ERDF A way of making Europe” and by MCIN/AEI/10.13039/501100011033. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project Nos. 2021/3-42, 2021/6-326, 2021-1-24 and 2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973. FS is partly supported by the Italian INFN program on Theoretical Astroparticle Physics (TAsP), by the French CNRS grant IEA “DaCo: Dark Connections” and by COST (European Cooperation in Science and Technology) via the COST Action COSMIC WISPers CA21106.
@@ Appendix A Uncertainties in the predicted profiles and mean distance travelled by positrons
In our simulations, we employ the gas maps developed by the GALPROP team [94, 95], which is one of the most popular gas distributions employed in CR propagation studies. A different gas distribution is not expected to change our conclusions significantly, but it may marginally affect our predictions for the zones very close to the GC, since different gas distributions can vary noticeably around the GC, often predicting a higher gas density than the one used in this work (albeit not significantly modifying the predictions shown). We note that the available gas distributions in current CR propagation codes are not optimized for very small scales (tens of parsecs) around the GC. Using a gas distribution with a higher density around the GC (which could happen if the region in the vicinity of Sgr A* concentrates more gas than assumed here) would lead to higher energy losses, which means that the positrons will remain closer to the DM distribution while thermalizing and, therefore, result in a more peaked emission profile and a higher bulge-to-disk ratio, which would further reinforce our conclusions.
@@
In terms of uncertainties related to the propagation parameters, it has already been pointed out that reacceleration is the main source of uncertainty for sub-GeV DM in Ref. [48]. We have tested how the line emission profile changes for two extreme benchmark cases of no reacceleration ( km/s) and very high reacceleration ( km/s). Different values for the halo height will only rescale our predictions, with no relevant implications on the shape of the profile. The difference in the predicted profile of the line is small, as can be seen in Fig. 8 (dashed lines) for the case of a MeV DM particle, following the Gondolo-Silk spike model. We notice that for higher masses, the effect of reacceleration is even lower.
@@
Second, as shown in Eq. (2.16), the shape of the line also depends on the distribution of ambient electrons available to form positronium bound states. Our benchmark predictions estimate a flat distribution of electrons along the Galactic disk and an exponential decrease away from it. We have tested how the assumption of a flat electron distribution in the Galactic disk changes our predictions. First, we have adopted the Ferriere distribution [96] of ionized hydrogen as implemented in the DRAGON2 code. The predicted profile is shown as a green line in Fig. 8. We note that the profile of the line is roughly identical to the one obtained in our benchmark case. To show an extreme case, we perform a similar test using the Nakanishi model, based on Ref. [97]. The predicted profile in this case is much more peaked, as can be seen from the blue line in Fig. 8. This, again, will simply decrease the fraction of DM contributing to the Galactic diffuse -ray emission in the regions where available measurements exist (the , region, for instance). We remark here on the need for measurements of the diffuse emission at smaller regions around the GC to probe these kinds of profiles. The last major ingredient that can affect the spatial distribution of the line here is how the different gas media are distributed and their temperature, which would change the factor in Eq. (2.16). However, a detailed modelling of the different media and their temperatures is well beyond the scope of this work.
@@
We estimate the average distance travelled by positrons in a neutral gas medium (considering that He constitutes around of the gas) as , where D is the diffusion coefficient that we employ in this work and is the loss-time given by the ionization energy losses. This is shown as blue lines in Fig. 9, where the solid lines represent the results for a cm-3 medium and the dashed lines the same but for a cm-3. We provide a rough estimate of the effect of reacceleration (green lines in Fig. 9) simply by evaluating the energy at which the positron spectra peak when including reacceleration (with our best-fit km/s) and assigning the average distance travelled by positrons of this energy. As one can see in Ref. [48], positrons injected by a MeV DM particle get an average energy of a few tens of MeV, and above MeV its effect becomes very minor. Since this is an averaged distance, we remark that of the injected positrons actually will travel longer distances before thermalizing. As a reference, a medium with density cm-3 will travel an average distance times lower than in the cm-3 case.
@@ Appendix B In-flight annihilation emission from sub-GeV DM
Finally, we also show the equivalent of Fig. 6 but for the NFW DM distribution in Fig.12. As we see, while the in-flight annihilation emission is similar to the other spike distributions, since it depends mostly on the mass of the DM particle, the predicted shape of the keV line distribution is unable to provide a good reproduction of SPI data. The corresponding values are cm3/s, cm3/s, cm3/s, cm3/s and cm3/s for DM masses of MeV, MeV, MeV, MeV and MeV, respectively.
@@ Appendix C Other regions of interest
This appendix is intended to compare the predicted in-flight positron annihilation emission with other publicly available observations of the diffuse MeV Galactic emission, especially by COMPTEL. In particular, the in-flight positron annihilation fluxes are calculated in a similar fashion to the right panels of Fig. 6 (i.e. taking the cross sections that provide the best-fit of the keV line emission to the longitude profile of SPI data). These comparisons are illustrated in Fig. 13, that show the predicted emissions for a Gondolo-Silk profile with COMPTEL and EGRET measurements of the Galactic diffuse emission in the - (top-left panel), - (top-right panel) and the - (bottom panels) region. In addition, the background components (bremsstrahlung and IC) are shown as dashed lines and the total background emission is shown as a solid black line.

### v3

@@ 3.2 Continuum emission
This is evidenced from Fig. 6, where we compare, in the right panels, the predicted in-flight annihilation emission for DM masses of to MeV, with SPI and COMPTEL data in the region of the sky, for a Gondolo-Silk (top row) and a *MIN (bottom row) DM spike benchmarks. For each DM mass, we perform a fit to the keV longitudinal profile (shown in the left panels) and normalize the value to the best-fit, including the astrophysical disk component from stars. The expected IC background emission is shown as a solid black line.
From Fig. 6, one can clearly see that masses of up to MeV seem to be well compatible with the current measurements of the diffuse -ray galactic flux, while higher masses will significantly exceed the data (especially, COMPTEL observations) and would exceed the background emission, which is not expected. A similar figure to this one is shown in Appendix A (Figure 12) but for the case of a NFW DM distribution without spike.
For completeness, we show in Table 1 our best-fit values used to obtain the estimations. We note that, as the DM mass increases, the best-fit cross section does not scale simply proportionally to with , as one would expect simply by the DM number density and as found indeed by previous fits that ignored positron propagation (see e.g. [64]). We instead find , that can be understood as follows: our signal is proportional to the number of thermalized positrons, and the larger the energy injection of the positrons (i.e. the DM mass) the less the amount of positrons that reach thermal energies before leaving the central regions of the Galaxy. Therefore larger DM masses need a slightly larger cross-section with respect to fits that ignore propagation. Finally on the cross-sections, building particle DM models that realize them goes beyond the purposes of this paper. Still, we note that these best-fit values are compatible with existing indirect detection limits (see e.g. [81] for a review), and that they can easily give rise to the correct DM relic abundance via thermal freeze-out, for example in the cases of -wave annihilations or of co-annihilations [24].
| 2 MeV | 5 MeV | 10 MeV | 20 MeV | 50 MeV | |
@@
A possible test of our proposal would come from data in ROIs other than -, as well as from different experiments. To the best of our knowledge, these are provided by COMPTEL and EGRET, which we show and discuss in App. C, finding that they do not alter our conclusions. We also include similar comparisons with SPI data in a ROI of around the GC, reaching energies of up to MeV (Fig. 16 in App. C). An important point of this dataset is that systematic uncertainties below keV (where IC on starlight becomes dominant) are better accounted for.
@@ 4 Summary and discussion
For the first time in this context, we have considered the possibility that the DM density profile develops a spike around Sgr A*. This is expected for peaked DM profiles like NFW, that were already known to be needed to possibly explain this signal. We have computed the associated 511 keV line emission and demonstrated that it can fit the bulge data whilst producing very little emission at the disk, which we have assumed to have originated mainly from astrophysical sources, e.g. young stars. Remarkably, this conclusion holds not only for Gondolo-Silk spikes, but also for more conservative ones that take into account DM profile softening by stars in the vicinity of Sgr A* and are supported by simulations (all our spike benchmarks respect the limits on the mass close to Sgr A* by the GRAVITY experiment [37]). DM profiles like NFW without a spike, that explained the observed 511 keV bulge emission, have been recently found to produce too much disk emission –which we confirmed here– and hence they were not favoured by data. Our findings, therefore, motivate new interest in the DM hypothesis for the 511 keV line. Although we remark that the measurements of the distribution of the line emission are very challenging and are subject to important systematic uncertainties, our work shows that spike profiles can reproduce the data satisfactorily adopting the positron propagation that is extrapolated from CR observations at higher energies.
@@
Finally concerning our new results, we have tested how some variations, either in the diffusion setup or in the electron density distribution in the Milky Way, can affect the reported spatial profiles of the emission. As we discuss in Appendix A, variations of our benchmark scenario mainly lead to a more peaked profile, which leads to an even higher ratio for the bulge-to-disk emission and thus reinforce our conclusions. Furthermore, an inhomogeneous diffusion scenario where CRs are more confined around the GC [82], as motivated by observations of the Fermi-LAT [83, 84, 85], will also lead to a more peaked profile of the predicted line. However, winds in the GC will have the opposite effect.
We now move to comment on observational implications of our results for MeV telescopes. First, we note that the SPI measurements employed are significantly affected by the templates used to extract the observations. In particular, the data that we are using are extracted using Gaussian templates, which have a similar symmetry as the one expected by DM-like models. Recently, stellar templates for the nuclear stellar bulge and a boxy bulge were shown to be statistically preferred over the profile expected from an NFW DM distribution [9], even though positron propagation was not included on top of NFW and thus, strictly speaking, the conclusion applies only to DM masses closest to an MeV. In addition, Ref. [9] found that the NFW template was the best model when analyzing the keV line, while it is the continuum o-ps emission what makes their analysis favour the stellar bulge hypothesis. Moreover, the templates of [9] did not consider the DM spike profiles that we are exploring here. Therefore, our goal here is not to provide the best and most accurate explanation of the keV line emission, which would require access to SPI’s raw data and the coded mask response function, but to demonstrate that the keV observations and the correlated in-flight annihilation emission motivate considering the possibility of an origin from DM annihilations alleviating previous constraints.
@@ Acknowledgements
We thank Michael Mancini and Paolo Panci for useful comments on the manuscript. SB and MF are supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union “NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, wich are both funded “ERDF A way of making Europe” and by MCIN/AEI/10.13039/501100011033. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project Nos. 2021/3-42, 2021/6-326, 2021-1-24 and 2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973. FS is partly supported by the Italian INFN program on Theoretical Astroparticle Physics (TAsP), by the French CNRS grant IEA “DaCo: Dark Connections” and by COST (European Cooperation in Science and Technology) via the COST Action COSMIC WISPers CA21106.
@@ Appendix A Uncertainties in the predicted profiles and mean distance travelled by positrons
In our simulations, we employ the gas maps developed by the GALPROP team [94, 95], which is one of the most popular gas distributions employed in CR propagation studies. A different gas distribution is not expected to change our conclusions significantly, but it may affect our predictions for the zones very close to the GC, since different gas distributions can vary noticeably around the GC, often predicting a higher gas density than the one used in this work (albeit not significantly modifying the predictions shown). We note that the available gas distributions in current CR propagation codes are not optimized for very small scales (tens of parsecs) around the GC. Using a gas distribution with a higher density around the GC (which could happen if the region in the vicinity of Sgr A* concentrates more gas than assumed here) would lead to higher energy losses, which means that the positrons will remain closer to the DM distribution while thermalizing and, therefore, result in a more peaked emission profile and a higher bulge-to-disk ratio, which would further reinforce our conclusions.
@@
In terms of uncertainties related to the propagation parameters, it has already been pointed out that reacceleration is the main source of uncertainty for sub-GeV DM in Ref. [48]. We have tested the effect on the spatial distribution of the line for two extreme benchmark cases of no reacceleration ( km/s) and very high reacceleration ( km/s). Different values for the halo height will only rescale our predictions, with no relevant implications on the shape of the profile. The difference in the predicted profile of the line is small, as can be seen in Fig. 8 (dashed lines) for the case of a MeV DM particle, following the Gondolo-Silk spike model. We notice that for higher masses, the effect of reacceleration is even lower.
@@
Second, as shown in Eq. (2.16), the spatial morphology of the line also depends on the distribution of ambient electrons available to form positronium bound states. Our benchmark predictions estimate a flat distribution of electrons along the Galactic disk and an exponential decrease away from it. We have tested how the assumption of a flat electron distribution in the Galactic disk changes our predictions. First, we have adopted the Ferriere distribution [60] of ionized hydrogen as implemented in the DRAGON2 code. The predicted profile is shown as a green line in Fig. 8. The predicted profile in this case is much more peaked, as can be seen from the green line in Fig. 8. To show a different case, we perform a similar test using the Nakanishi model, based on Ref. [59]. Finally, as explained in the text, we also tested the convolution with a distribution that results from an even combination of the Ferriere and Nakanishi models, as a blue line (“Combined dist” in the legend). This, again, will simply decrease the fraction of DM contributing to the Galactic diffuse -ray emission in the regions where available measurements exist (the , region, for instance). We remark here on the need for measurements of the diffuse emission at smaller regions around the GC to probe these kinds of profiles. The last major ingredient that can affect the spatial distribution of the line here is how the different gas media are distributed and their temperature, which would change the factor in Eq. (2.16). However, a detailed modeling of the different media and their temperatures is well beyond the scope of this work.
@@
We estimate the average distance traveled by positrons in a neutral gas medium (considering that He constitutes around of the gas) as , where D is the diffusion coefficient that we employ in this work and is the loss-time given by the ionization energy losses. This is shown as blue lines in Fig. 9, where the solid lines represent the results for a cm-3 medium and the dashed lines the same but for a cm-3. We provide a rough estimate of the effect of reacceleration (green lines in Fig. 9) simply by evaluating the energy at which the positron spectra peak when including reacceleration (with our best-fit km/s) and assigning the average distance travelled by positrons of this energy. As one can see in Ref. [48], positrons injected by a MeV DM particle get boosted and their spectrum peaks at an average energy of a few tens of MeV, while for DM masses above MeV its effect becomes very minor. Since this is an averaged distance, we remark that of the injected positrons actually will travel longer distances before thermalizing. As a reference, a medium with density cm-3 will travel an average distance times lower than in the cm-3 case.
@@ Appendix B In-flight annihilation emission from sub-GeV DM
Finally, we also show the equivalent of Fig. 6 but for the NFW DM distribution in Fig.12. As we see, while the in-flight annihilation emission is similar to the other spike distributions, since it depends mostly on the mass of the DM particle, the predicted spatial morphology of the keV line distribution is unable to provide a good reproduction of SPI data. The corresponding values are cm3/s, cm3/s, cm3/s, cm3/s and cm3/s for DM masses of MeV, MeV, MeV, MeV and MeV, respectively.
@@ Appendix C Other regions of interest
This appendix is intended to compare the predicted in-flight positron annihilation emission with other publicly available observations of the diffuse MeV Galactic emission, especially by COMPTEL.
First, we show the maps of the expected in-flight annihilation emission in the inner , in Fig. 13, for a 20 MeV DM mass and cross sections of cm3/s and cm3/s, for the Gondolo-Silk profile (left panel) and *Min profile (right panel), respectively. As we can see, the effect of the exponential suppression of the electron density with latitude is evident in both cases.
Then, we focus on the comparisons of the expected in-flight positron annihilation emission signals for other ROIs. In these comparisons, the in-flight positron annihilation fluxes are calculated in a similar fashion to the right panels of Fig. 6 (i.e. taking the cross sections that provide the best-fit of the keV line emission to the longitude profile of SPI data). These comparisons are illustrated in Fig. 14, that show the predicted emissions for a Gondolo-Silk profile with COMPTEL and EGRET measurements of the Galactic diffuse emission in the - (top-left panel), - (top-right panel) and the - (bottom panels) region. In addition, the background components (bremsstrahlung and IC) are shown as dashed lines and the total background emission is shown as a solid black line.
@@
Finally, we also show in Fig. 16 similar comparisons but for the SPI measurements in the around the GC. As it is expected, the contribution to the 511 keV line emission and the continuum emission is much lower than in more central regions of interest.